How to prepare 500ml of 50percent of ethanol from a bag of 90 percent of ethanol and water.
Answer by stanbon(75887)

How to prepare 500ml of 50percent of ethanol from a bag of 90 percent of ethanol and water.
-----------------
Equation:
water + water = water
0.10x + 1*(500-x) = 0.50*500
--------------
Multiply thru by 100 to get:
10x + 100*500 - 100x = 50*500
-90x = -50*500
x = (5/9)500
x = 277.78 ml
500-x = 222.22 ml (amt of pure water to add)
===============
Cheers,
Stan H.
---------------